A new product (Brand B) will be introduced. This new product has certain advantages that were not found in the firm’s existing products. Due to the higher costs, however, this new product will carry new a new contribution margin of $10 per unit (Compared with the margin of $12 per unit from the existing product-Brand A). If the firm plans to sell 500 units of the new product (Brand B) in the first year with the 90% (Unit cannibalization rate). If the firm plans to sell 500 units of the “Brand B” in its first year, please calculate the overall contribution. Should this firm introduce this new product in the first year
Answer #1

Brand A contribution = $ 12 per unit

Brand B contribution = $ 10 per unit

Sale of brand B = 500 units

Cannibalization = 90%

500 unit of Brand B = 90% of sale of brand A

Sale of brand A before cannibalization = 500 * 100/90 = 555.55 ~ 556 units

Sale of brand A after cannibalization = 556 - 500 = 56 units

total contribution = contribution from brand A + contribution from brand B

= (56 * 12) + (500 * 10) = $ 5672

Total contribution before launch of brand B = 556 * 12 = $ 6672

Loss in contribution by launching brand B = 6672 - 5672 = $ 1000

Therefore the firm should not introduce the product in the first year.